Red gold is a gold‑copper alloy used to make jewelry. A piece of jewelry made of red gold weighs 8.80 g and has a volume of 0.594 cm3. Gold has a density of 19.3 g/cm3 and copper has a density of 8.96 g/cm3

. Calculate the percentage by mass of each metal in the jewelry. Assume the total volume of the jewelry is the sum of the volumes of the two metals it contains.

gold:

%

copper:

%

Pure gold is defined as having 24 carats. When mixed in an alloy, the carats of gold are given as a percentage of this value. For example, a piece of jewelry made with 50% gold has 12 carats. State the purity of this piece of red gold jewelry in carats.

purity:

carats
Answer #1

solution - In this question, we have - weight of red Gold Jewelry = 8.80g - volume of Jewelry = 0.594 cm3, Density of Gold =So, volume of gold used = 4 cm² - 0.337 cm3 volume of copper used = y cm3 = (0.594-2) (m3 = (0.594 -0.337) cm = 0,257 cm - masimilarly, mass % of copper :- » mass t of copper = 2,3 g x 100% 8.89 mass % of copper = 26.1% If jewelry is made up of 100%